OpenStudy (anonymous):

A sign in a clothing store said, "All prices reduced 25%. Final prices marked on tags." DeLea found a sweatshirt that showed a price of $27. What was the original price of the sweatshirt?

OpenStudy (anonymous):

75% of the old price is the new price. Let x be the old price. Solve the following for x: 0.75 x = 27 x = 27/0.75 = $36 Checking the answer:\[\frac{36-27}{36}100= 25\% \text{ off}\]

OpenStudy (anonymous):

Cool